MILLIMETER-WAVE SPECTROSCOPY OF THE FeCO(X^{3}Σ^{−}) AND FeNO(X^{2}Δ_{i}) RADICALS IN THE VIBRATIONAL EXCITED STATES

\underline{SEIKI IKEDA}, MOTOKI NAKASHIMA, AI SAKAMOTO, MASATO HAYASHI, KEUSUKE HARADA AND KEIICHI TANAKA, Department of Chemistry, Faculty of Science, Kyushu University, Hakozaki, Higashiku, Fukuoka, 812-8581 JAPAN.

Rotational spectra of the FeCO and FeNO radicals produced by UV laser photolysis in the vibrational excited states were measured in the millimeter-wave region with the conventional absorption cell at room temperature.

The rotational transitions of the FeCO radical in the ground and ν_{2} states have been observed by millimeter-wave spectroscopy^{a}, and the ν_{1} fundamental band and hot band from ν_{2} state also have been studied by infrared diode laser spectroscopy. In the present work, the rotational transitions (J = 33 - 32 ∼ 37 - 36 ) in the ν_{3} state of the X^{3}Σ^{−} state were observed to split into 3 components due to the spin-rotation and spin-spin interactions. Molecular constants including rotational constant and centrifugal distortion constant were determined by a least squares fitting. The equilibrium rotational constant B_{e} was calculated to be 4373.405(72) MHz from the vibration rotation constant α_{3} = 20.2051(42) MHz, and previously reported α_{1} and α_{2}. The bond length between Fe and C, calculated to be 1.725 Å assuming r_{CO}= 1.159 Å, agrees well with the ab initio result, r_{FeC} = 1.722 Å^{b}. The 2ν_{2} state split into 9 substates due to the vibronic interaction, and the rotational transitions in the P = 0 component were observed.

The rotational transitions of the FeNO radical in the ground and ν_{2} states^{c}, and the ν_{1} band have been observed in the millimeter-wave and infrared region, respectively. The rotational transitions (J = 28.5 - 27.5 ∼ 32.5 - 31.5 ) in the 2ν_{2} state of the X^{2}Δ_{i} state were observed in the present study. The 2ν_{2} state (Ω = 5/2) splits into 3 substates, 2Γ_{P =9/2},^{2} Δ_{P =5/2} and ^{2}Σ_{P =1/2}, due to the vibronic interaction. The absorption lines in the ^{2}Σ_{P =1/2} state split into two components because of the p-type doubling. The transition in the ν_{3} state is now under survey to determine the constant α_{3} and the equilibrium rotational constant.

aK. Tanaka, M. Shirasaka, T. Tanaka, J. Chem. Phys., 106, 6820-6824 (1997)

bU. Nagashima, R. Okuda, T. Hirano, Molecular Spectroscopy Symposium, Tokyo (2005)

cS. Ikeda, M. Nakashima, M. Hayashi, K. Harada and K. Tanaka, 60th Ohio Meeting (2005)